A solution is prepared by dissolving 19.2 g of glucose in 205 g of water. The final
volume of the solution is 214 ml. Find the concentration of the solution in units of
molality.
1
Expert's answer
2021-09-17T02:29:35-0400

Molarity = moles of the solute/litres of the solution

Molar mass of glucose = 180.156

= 19.2/180.156 = 0.107mol

214/1000 =0.214kg

= 0.107/0.214

= 0.5mol/Kg